# Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu

### v1

@@ 3.1.2 Parameter-Tuning Paradigm
Full-parameter instruction tuning has been shown to be an effective approach to bridge the gap between task-agnostic pre-training and task-specific inference. Specifically, Wang et al. (2022) design unified sentiment instruction for various aspect-based sentiment analysis tasks to elicit the LLMs. Varia et al. (2022) utilize task-specific sentiment instructions to fine-tune LLMs for the inter-task dependency. Yang and Li (2023) transform the visual input into plain text during prompt construction for instruction tuning. These works demonstrate the potential of tuning LLMs for advanced sentiment analysis.
@@
Sentiment analysis techniques have numerous real-world applications such as opinion mining Zhao et al. (2016). Therefore, efficiency is a vital dimension for evaluating sentiment analysis methods. Qiu et al. (2023) utilize LoRA to tune LLMs on the empathy multi-turn conversation dataset namely SMILECHAT to develop emotional support systems.
@@ 3.2 Information Extraction
Information Extraction (IE) tasks aim at extracting structural information from plain text, which typically includes relation extraction (RE), named entity recognition (NER), and event extraction (EE) Xu et al. (2023a).
@@ 3.2.1 Parameter-Frozen Paradigm
Inspired by the impressive capabilities of LLMs on various tasks, recent studies Zhang et al. (2023c); Wei et al. (2023a) begin to explore zero-shot prompting methods to solve IE tasks by leveraging knowledge embedded in LLMs. Wei et al. (2023a) , Xie et al. (2023) and Zhang et al. (2023c) propose a series of methods to decompose question-answering tasks by breaking down NER into smaller, simpler subproblems, which improves the overall process. In addition, Xie et al. (2023) further introduce two methods, syntactic prompting and tool augmentation, to improve LLMs’ performance by incorporating the syntactic information.
@@
Considering the gap between sequence labeling and text generation, providing exemplars could help LLMs better understand the given task and follow the problem-solving steps. To select pertinent demonstrations, Li and Zhang (2023) deploy the retrieval module to retrieve the most suitable examples for the given test sentence. Instead of using natural language for structured output, Li et al. (2023e) and Bi et al. (2023) propose reformulating IE tasks as code with code-related LLMs such as Codex.
@@ 3.2.2 Parameter-Tuning Paradigm
A common practice to customize LLMs is fine-tuning LLMs on the collected dataset. There typically are three tuning paradigms adopted to enhance LLMs’ abilities. The first one is tuning LLMs on a single dataset to strengthen a specific ability. The second one is standardizing data formats across all IE subtasks, thus enabling a single model to efficiently handle diverse tasks (Lu et al., 2023a; Gan et al., 2023). The last one is tuning LLMs on a mixed dataset and testing on the unseen tasks Sainz et al. (2023); Wang et al. (2023f), which is always used to improve the generalization ability of LLMs.
@@
Tuning huge parameters of LLMs poses a significant challenge to both research and development. To address this challenge, Das et al. (2023b) propose a method for dynamic sparse fine-tuning that focuses on a specific subset of parameters during the IE training process. This approach is particularly useful when dealing with limited data. Meanwhile, Liang et al. (2023) introduce Lottery Prompt Tuning (LPT), a method that efficiently tunes only a portion of the prompt vectors used for lifelong information extraction. This technique optimizes both parameter efficiency and deployment efficiency.
@@ 3.3 Dialogue Understanding
Dialogue understanding typically consists of spoken language understanding (SLU) Tur and De Mori (2011); Qin et al. (2019, 2021) and dialogue state tracking (DST) Sarikaya et al. (2016); Jacqmin et al. (2022).
@@ 3.3.1 Parameter-Frozen Paradigm
Recent studies highlight the effectiveness of LLMs in dialogue understanding through zero-shot prompting Pan et al. (2023); He and Garner (2023); Hudeček and Dušek (2023); Heck et al. (2023). Gao et al. (2023a) and Addlesee et al. (2023) introduce zero-shot chain-of-thought prompting strategies in LLMs, enhancing understanding by step-by-step reasoning. Moreover, Zhang et al. (2023i) and Wu et al. (2023c) treat SLU and DST as agent systems and code generation tasks to effectively improve task performance. Further, Chung et al. (2023), Chi et al. (2023) and Zhang et al. (2023h) extend the task to actual scenarios and understand the dialog by zero-shot prompting for efficient interaction and dialog management.
@@
Limited by the instruction following ability of the LLMs, recent studies have focused on improving model performance in dialogue understanding through the relevant few-shot demonstrations Hudeček and Dušek (2023). To address “overfitting” in the given few-shot demonstrations, Hu et al. (2022b), King and Flanigan (2023), Das et al. (2023a), Li et al. (2022b), Lee et al. (2023), King and Flanigan (2023) and Addlesee et al. (2023) further introduce some methods for retrieving diverse few-shot demonstrations to improve understanding performance. Lin et al. (2023) and Cao (2023) integrate DST tasks with an agent through in-context-learning, enhancing dialogue understanding capabilities.
@@ 3.3.2 Parameter-Tuning Paradigm
Full-parameter tuning involves not freezing any parameters and using all parameters to train dialogue understanding tasks Yu et al. (2022). Specifically, Xie et al. (2022); Zhao et al. (2022a) unifies structured tasks into a textual format by training full parameters demonstrating significant improvement and generalization. Gupta et al. (2022) utilize input with some demonstrations as a new DST representation format to train LLM with full parameters and achieve great results.
@@
Limited by the huge cost of full-parameter fine-tuning, a lot of work begins to focus more on Parameter-Efficient Tuning (PET) for lower-cost dialogue understanding task training. Specifically, Feng et al. (2023b) present LDST, a LLaMA-driven DST framework that leverages LoRA technology for parameter-efficient fine-tuning, achieving performance comparable to ChatGPT. Liu et al. (2023b) provide a key-value pair soft-prompt pool, selecting soft-prompts from the prompting pool based on the conversation history for better PET.
@@ 3.4 Table Understanding
Table understanding involves the comprehension and analysis of structured data presented in tables, focusing on interpreting and extracting meaningful information, like Table Question Answering Jin et al. (2022).
@@ 3.4.1 Parameter-Frozen Paradigm
Recently, the advancements for LLMs have paved the way for exploring zero-shot learning capabilities in understanding and interpreting tabular data Singha et al. (2023); Patnaik et al. (2024); Ye et al. (2024). Ye et al. (2023) and Sui et al. (2023a) concentrate on breaking down large tables into smaller segments to reduce irrelevant data interference during table understanding. Further, Patnaik et al. (2024) introduce CABINET, a framework that includes a module for generating parsing statements to emphasize the data related to a given question. Sui et al. (2023b) develop TAP4LLM, enhancing LLMs’ table understanding abilities by incorporating reliable information from external knowledge sources into prompts. Additionally, Ye et al. (2024) propose a DataFrameQA framework to utilize secure Pandas queries to address issues of data leakage in table understanding. These efforts signify a significant stride towards leveraging LLMs for more effective and efficient zero-shot learning in table data comprehension.
@@
Few-shot learning has been an increasingly focal point for researchers to address the limitations of LLMs, particularly in the context of table understanding and instruction following ability Chen (2023); Zhang et al. (2024). Luo et al. (2023b) propose a hybrid prompt strategy coupled with a retrieval-of-thought to further improve the example quality for table understanding tasks. Cheng et al. (2022) introduce Binder to redefine the table understanding task as a coding task, enabling the execution of code to derive answers directly from tables. Furthermore, Li et al. (2023b), Jiang et al. (2023) and Zhang et al. (2023k, f) conceptualize the table understanding as a more complex agent task, which utilizes external tools to augment LLMs in table tasks. Building upon these developments, ReAcTable Zhang et al. (2023j) integrates additional actions into the process, such as generating SQL queries, producing Python code, and directly answering questions, thereby further enriching the few-shot learning landscape for LLMs.
@@ 3.4.2 Parameter-Tuning Paradigm
Leveraging the existing capabilities of LLMs, Full-Parameter Tuning optimizes these models for specific table understanding tasks. Li et al. (2023d) and Xie et al. (2022) adapt a substantial volume of table-related data for table instruction tuning, which leads to better generalization in table understanding tasks. Additionally, Xue et al. (2023) introduce DB-GPT to enhance LLMs by fine-tuning them and integrating a retrieval-augmented generation component to better support table understanding.
@@
Xie et al. (2022) utilize prompt-tuning for efficient fine-tuning within a unified framework of table representation instructions. Moreover, Zhang et al. (2023a), Zhu et al. (2024) and Bai et al. (2023) adapt Low-Rank Adaptation (LoRA) during instruction-tuning for better table understanding and further table cleaning. Furthermore, Zhang et al. (2023d) address challenges related to long table inputs by implementing LongLoRA, demonstrating its efficacy in managing long-context issues in table understanding tasks.
@@ 4 Natural Language Generation
This section presents the LLMs for classific NLP generation tasks containing Summarization (§$\S$4.1), Code Generation (§$\S$4.2), Machine Translation (§$\S$4.3), and Mathematical Reasoning (§$\S$4.4), which are illustrated in Figure 3.
@@ 4.1 Summarization
Summarization aims to distill the most essential information from a text document, producing a concise and coherent synopsis that retains the original content’s primary themes (Shi et al., 2018).
@@ 4.1.1 Parameter-Frozen Paradigm
In the exploration of zero-shot learning for text summarization, LLMs such as GPT-3 have demonstrated amazing and superior performance in generating concise and factually accurate summaries, challenging the need for traditional fine-tuning approaches (Goyal et al., 2022; Bhaskar et al., 2022; Wang et al., 2023b). Zhang et al. (2023e) highlight instruction tuning as pivotal for LLMs’ summarization success. Ravaut et al. (2023b) scrutinize LLMs’ context utilization, identifying a bias towards initial document segments in summarization tasks. These studies collectively underscore the versatility and challenges of deploying LLMs in zero-shot summarization.
@@
For few-shot learning, LLMs like ChatGPT are scrutinized for their summarization abilities. Zhang et al. (2023b) and Tang et al. (2023) demonstrate that leveraging in-context learning and a dialog-like approach can enhance LLMs’ extractive summarization, particularly in achieving summary faithfulness. Adams et al. (2023) introduce a “Chain of Density” prompting technique, revealing a preference for denser, entity-rich summaries over sparser ones. Together, these studies reveal the evolving strategies to optimize LLMs for summarization tasks.
@@ 4.1.2 Parameter-Tuning Paradigm
Full-Parameter Tuning for text summarization leverages the power of LLMs, optimizing them for specific summarization tasks. DIONYSUS (Li et al., 2022a) adapts to new domains through a novel pre-training strategy tailored for dialogue summarization. Socratic Pretraining (Pagnoni et al., 2022) introduces a question-driven approach to improve the summarization process. This allows the model to be easily adapted for different summarization tasks, resulting in more controllable and relevant summaries.
@@
PET strategies have revolutionized the adaptability of large pre-trained models for specific summarization tasks, demonstrating the power of fine-tuning with minimal parameter adjustments Feng et al. (2023a). Zhao et al. (2022b) and Yuan et al. (2022) adapt prefix-tuning (Li and Liang, 2021b) for dialogue summarization, enhancing model knowledge and generalization across domains. Ravaut et al. (2023a) develop PromptSum to combine prompt tuning with discrete entity prompts for controllable abstractive summarization. These approaches collectively show the efficacy of PET in enabling robust, domain-adaptive, and controllable summarization with minimal additional computational costs.

### v2

@@ 3.1.2 Parameter-Tuning Paradigm
Full-parameter instruction tuning has been shown to be an effective approach to bridge the gap between task-agnostic pre-training and task-specific inference. Specifically, Wang et al. [120] design unified sentiment instruction for various aspect-based sentiment analysis tasks to elicit the LLMs. Varia et al. [121] utilize task-specific sentiment instructions to fine-tune LLMs for the inter-task dependency. Yang and Li [122] transform the visual input into plain text during prompt construction for instruction tuning. Moreover, Zhang et al. [47] conduct an empirical study to evaluate bLLMs’ effectiveness in sentiment analysis for software engineering, revealing their advantages in data-scarce scenarios and limitations compared to fine-tuned sLLMs with sufficient training data. These works demonstrate the potential of tuning LLMs for advanced sentiment analysis.
@@
Sentiment analysis techniques have numerous real-world applications such as opinion mining [123]. Therefore, efficiency is a vital dimension for evaluating sentiment analysis methods. Qiu et al. [124] utilize LoRA to tune LLMs on the empathy multi-turn conversation dataset namely SMILECHAT to develop emotional systems.
@@ 3.2 Information Extraction
Information Extraction (IE) tasks aim at extracting structural information from plain text, which typically includes relation extraction (RE), named entity recognition (NER), and event extraction (EE) [182].
@@ 3.2.1 Parameter-Frozen Paradigm
Inspired by the impressive capabilities of LLMs on various tasks, recent studies [48, 24] begin to explore zero-shot prompting methods to solve IE tasks by leveraging knowledge embedded in LLMs. Wei et al. [24] , Xie et al. [49] and Zhang et al. [48] propose a series of methods to decompose question-answering tasks by breaking down NER into smaller, simpler subproblems, which improves the overall process. Xie et al. [49] introduce two methods, syntactic prompting and tool augmentation, to improve performance of LLMs by incorporating the syntactic information. Siepmann et al. [183] explore the use of GPT-4 for automated information extraction of diagnoses, medications, and allergies from discharge letters, demonstrating high accuracy with prompt tuning and highlighting its potential to reduce administrative burden in healthcare. In addition, Gu et al. [184] conduct a cross-sectional study to evaluate advanced open-source LLMs for information extraction of social determinants of health (SDoH) from clinical notes, using human-annotated EHR data and comparing against a pattern-matching baseline.
@@
Considering the gap between sequence labeling and text generation, providing exemplars could help LLMs better understand the given task and follow the problem-solving steps, especially in tasks requiring structured outputs and clear format adherence for accuracy. To select pertinent demonstrations, Li and Zhang [50] deploy the retrieval module to retrieve the most suitable examples for the given test sentence, aiming to enhance task relevance and response accuracy. Instead of using natural language for structured output, Li et al. [51] and Bi et al. [52] propose reformulating IE tasks as code with code-related LLMs such as Codex, effectively leveraging their powerful syntax-aware generation and reasoning capabilities. Fornasiere et al. [53] introduce prompt-based strategies for small-scale LLMs to extract structured and unstructured medical information from clinical texts, demonstrating strong zero-shot performance and enhanced explainability through line-number referencing to source text. Tang et al. [54] explore the impact of various prompt engineering strategies, persona, chain-of-thought, and few-shot prompting, on the performance of GPT-3.5 and GPT-4 in extracting key information from medical publications, evaluating alignment with ground truth using multiple comprehensive metrics.
@@ 3.2.2 Parameter-Tuning Paradigm
A common practice to customize LLMs is fine-tuning LLMs on the collected dataset. There typically are three tuning paradigms adopted to enhance LLMs’ abilities. The first one is tuning LLMs on a single dataset to strengthen a specific ability. The second one is standardizing data formats across all IE subtasks, thus enabling a single model to efficiently handle diverse tasks [125, 126]. The last one is tuning LLMs on a mixed dataset and testing on the unseen tasks [127, 128], which is always used to improve the generalization ability of LLMs. Rixewa et al. [133] introduce a unified interleaved representation with cross-modal attention to enhance multi-modal information retrieval, enabling accurate and efficient processing of complex content across text and image formats.
@@
Tuning huge parameters of LLMs poses a significant challenge to both research and development. To address this challenge [185, 186], Das et al. [129] propose a method for dynamic sparse fine-tuning that focuses on a specific subset of parameters during the IE training process. This approach is particularly useful when dealing with limited data. Meanwhile, Liang et al. [130] introduce Lottery Prompt Tuning (LPT), a method that efficiently tunes only a portion of the prompt vectors used for lifelong information extraction. This technique optimizes both parameter efficiency and deployment efficiency. Dagdelen et al. [131] introduce a simple and flexible approach to fine-tuning LLMs for joint named entity recognition and relation extraction, enabling the generation of structured scientific knowledge records from complex materials chemistry texts. Xue et al. [132] introduce AutoRE, a novel end-to-end document-level relation extraction model using the RHF paradigm and parameter-efficient fine-tuning, enabling state-of-the-art performance without relying on predefined options.
@@ 3.3 Dialogue Understanding
Dialogue understanding typically consists of spoken language understanding (SLU) [187] and dialogue state tracking (DST) [188].
@@ 3.3.1 Parameter-Frozen Paradigm
Recent studies highlight the effectiveness of LLMs in dialogue understanding through zero-shot prompting [55, 56, 57, 58, 189]. Gao et al. [59] and Addlesee et al. [68] introduce zero-shot chain-of-thought prompting strategies in LLMs, enhancing understanding by step-by-step reasoning. Moreover, Zhang et al. [61] and Wu et al. [63] treat SLU and DST as agent systems and code generation tasks to effectively improve task performance. Further, Chung et al. [69], Chi et al. [65] and Zhang et al. [62] extend the task to actual scenarios and understand the dialog by zero-shot prompting for efficient interaction and dialog management. Recently, Qin et al. [190] and Qin et al. [191] propose a series of multi-stage solution frameworks that leverages the interactive capabilities of LLMs to address single‑intent and multi‑intent SLU tasks respectively. Dong et al. [192] propose a multi-agent framework, ProTOD, which is a novel active DST planner framework based on multiple LLMs’ interation, designed to enhance the dialog’s proactivity and goal completion rate.
@@
Limited by the instruction following ability of the LLMs, recent studies have focused on improving model performance in dialogue understanding through the relevant few-shot demonstrations [57]. To address “overfitting” in the given few-shot demonstrations, Hu et al. [66], King and Flanigan [67], Das et al. [64], Li et al. [60], Lee et al. [70], King and Flanigan [67] and Addlesee et al. [68] further introduce some methods for retrieving diverse few-shot demonstrations to improve understanding performance. Lin et al. [71] and Cao [72] integrate DST tasks with an agent through in-context-learning, enhancing dialogue understanding capabilities.
@@ 3.3.2 Parameter-Tuning Paradigm
Full-parameter tuning involves not freezing any parameters and using all parameters to train dialogue understanding tasks [137]. Specifically, Xie et al. [134], Zhao et al. [135] unifies structured tasks into a textual format by training full parameters demonstrating significant improvement and generalization. Gupta et al. [136] utilize input with some demonstrations as a new DST representation format to train LLM with full parameters and achieve great results. Acikgoz et al. [193] suggest that DST, typically trained on a limited set of APIs, needs new data for quality maintenance. They propose a unified instruction-tuning paradigm for multi-turn DST and advanced function calls, enhancing dialogue management and generalization.
@@
Limited by the huge cost of full-parameter fine-tuning, a lot of work begins to focus more on Parameter-Efficient Tuning (PET) for lower-cost dialogue understanding task training. Specifically, Feng et al. [138] present LDST, a LLaMA-driven DST framework that leverages LoRA technology for parameter-efficient fine-tuning, achieving performance comparable to ChatGPT. Liu et al. [139] provide a key-value pair soft-prompt pool, selecting soft-prompts from the prompting pool based on the conversation history for better PET. Further Yin et al. [194] address the multi-intent detection task and introduces MIDLM, a bidirectional LLM framework that enables autoregressive LLMs to leverage bidirectional information through post-training, thereby eliminating the need to train the model from scratch.
@@ 3.4 Table Understanding
Table understanding involves the comprehension and analysis of structured data presented in tables, focusing on interpreting and extracting meaningful information, like Table Question Answering [195, 196, 197].
@@ 3.4.1 Parameter-Frozen Paradigm
Recently, the advancements for LLMs have paved the way for exploring zero-shot learning capabilities in understanding and interpreting tabular data [73, 74, 76]. Ye et al. [75] and Sui et al. [77] concentrate on breaking down large tables into smaller segments to reduce irrelevant data interference during table understanding. Further, Patnaik et al. [74] introduce CABINET, a framework that includes a module for generating parsing statements to emphasize the data related to a given question. Sui et al. [78] develop TAP4LLM, enhancing LLMs’ table understanding abilities by incorporating reliable information from external knowledge sources into prompts. Additionally, Ye et al. [76] propose a DataFrameQA framework to utilize secure Pandas queries to address issues of data leakage in table understanding. These efforts signify a significant stride towards leveraging LLMs for more effective and efficient zero-shot learning in table data comprehension.
@@
Few-shot learning has been an increasingly focal point for researchers to address the limitations of LLMs, particularly in the context of table understanding and instruction following ability [84, 83, 198]. Luo et al. [85] propose a hybrid prompt strategy coupled with a retrieval-of-thought to further improve the example quality for table understanding tasks. Cheng et al. [79] introduce Binder to redefine the table understanding task as a coding task, enabling the execution of code to derive answers directly from tables. Furthermore, Li et al. [86], Jiang et al. [87] and Zhang et al. [81, 80] conceptualize the table understanding as a more complex agent task, which utilizes external tools to augment LLMs in table tasks. Building upon these developments, ReAcTable [82] integrates additional actions into the process, such as generating SQL queries, producing Python code, and directly answering questions, thereby further enriching the few-shot learning landscape. Wang et al. [88] introduce Chain-of-Table, a framework that guides LLMs to perform table-based reasoning by iteratively updating tabular data as intermediate steps, enabling structured, dynamic reasoning chains that significantly improve performance on table understanding tasks. Kong et al. [89] propose OpenTab, an open-domain table reasoning framework that enhances LLMs’ ability to handle structured table data by retrieving relevant tables and generating SQL programs for reasoning, significantly improving accuracy over existing methods in both open and closed domain scenarios.
@@ 3.4.2 Parameter-Tuning Paradigm
Leveraging the existing capabilities of LLMs, Full-Parameter Tuning optimizes these models for specific table understanding tasks. Li et al. [140] and Xie et al. [134] adapt a substantial volume of table-related data for table instruction tuning, which leads to better generalization in table understanding tasks. Additionally, Xue et al. [141] introduce DB-GPT to enhance LLMs by fine-tuning them and integrating a retrieval-augmented generation component to better support table understanding.
@@
Xie et al. [134] utilize prompt-tuning for efficient fine-tuning within a unified framework of table representation instructions. Moreover, Zhang et al. [142], Zhu et al. [143] and Bai et al. [144] adapt Low-Rank Adaptation (LoRA) during instruction-tuning for better table understanding and further table cleaning. Furthermore, Zhang et al. [145] address challenges related to long table inputs by implementing LongLoRA, demonstrating its efficacy in managing long-context issues in table understanding tasks. He et al. [146] introduce TableLoRA, a table-specific fine-tuning module that enhances LLMs’ understanding of tabular data under parameter-efficient settings by combining specialized table serialization and 2D positional encoding to improve performance on structured table tasks. Li et al. [147] introduce a new “table fine-tuning” paradigm that enhances language models like GPT-3.5 and ChatGPT on diverse table-understanding tasks by training them with synthesized table-based instructions, significantly improving their performance and generalizability on structured tabular data.
@@ 4 Natural Language Generation
This section presents the LLMs for classific NLP generation tasks containing Summarization ($\S$4.1), Code Generation ($\S$4.2), Machine Translation ($\S$4.3), and Mathematical Reasoning ($\S$4.4), which are illustrated in Figure 3.
@@ 4.1 Summarization
Summarization aims to distill the essential information from a text document, producing a concise and coherent synopsis that retains the original content’s themes [199].
@@ 4.1.1 Parameter-Frozen Paradigm
In the exploration of zero-shot learning for text summarization, LLMs such as GPT-3 have demonstrated amazing and superior performance in generating concise and factually accurate summaries, challenging the need for traditional fine-tuning approaches [90, 92, 93]. Zhang et al. [94] highlight instruction tuning as pivotal for LLMs’ summarization success. Ravaut et al. [91] scrutinize LLMs’ context utilization, identifying a bias towards initial document segments in summarization tasks [200, 201]. Furthermore, Yun et al. [202] enhances automatic summarization by integrating human interaction and semantic graphs, enabling the generation of higher-quality, personalized summaries tailored to individual users’ interests and needs. These studies collectively underscore the versatility and challenges of deploying LLMs in zero-shot summarization.
@@
For few-shot learning, LLMs like ChatGPT are scrutinized for their summarization abilities. Zhang et al. [95] and Tang et al. [97] demonstrate that leveraging in-context learning and a dialog-like approach can enhance LLMs’ extractive summarization, particularly in achieving summary faithfulness. Adams et al. [96] introduce a “Chain of Density” prompting technique, revealing a preference for denser, entity-rich summaries over sparser ones. Moreover, recent studies have begun to leverage the reflective capabilities [203], deeper reasoning abilities [204], and planning abilities [205] of large reasoning models to enhance the depth of thought as well as the conciseness and clarity of summaries. Together, these studies reveal the evolving strategies to optimize LLMs for summarization tasks.
@@ 4.1.2 Parameter-Tuning Paradigm
Full-Parameter Tuning for text summarization leverages the power of LLMs, optimizing them for specific summarization tasks. DIONYSUS [206] adapts to new domains through a novel pre-training strategy tailored for dialogue summarization. Socratic Pretraining [148] introduces a question-driven approach to improve the summarization process. Further, Wang et al. [207] and Lu et al. [208] demonstrate that carefully prompting LLMs produces well‑structured rationales, which can guide smaller models with fully tuning to generate summaries that are both more concise and of higher quality. More recently, Aali et al. [209] and Wu et al. [210] employ meticulously annotated supervised fine-tuning (SFT) data and prediction feedback-based reinforcement learning, respectively, enabling their models to match or even surpass the performance of proprietary closed-source models. Overall, this allows the model to be easily adapted for different summarization tasks, resulting in more controllable and relevant summaries.
@@
PET strategies have revolutionized the adaptability of large pre-trained models for specific summarization tasks, demonstrating the power of fine-tuning with minimal parameter adjustments [151]. Zhao et al. [149] and Yuan et al. [150] adapt prefix-tuning [152] for dialogue summarization, enhancing model knowledge and generalization across domains. Ravaut et al. [153] develop PromptSum to combine prompt tuning with discrete entity prompts for controllable abstractive summarization. These approaches collectively show the efficacy of PET in enabling robust, domain-adaptive, and controllable summarization with minimal additional computational costs.